OpenStudy (anonymous):

need help with this please! Integrate and write a non-trivial equation 10^(x+y)

OpenStudy (anonymous):

*non trivial solution

OpenStudy (asnaseer):

are you familiar with wolfram alpha web site?

OpenStudy (asnaseer):

if so, you can enter your equation there and will be tell you what the integral is including the steps taken to get it. http://www2.wolframalpha.com/input/?i=10%5E%28x%2By%29

OpenStudy (asnaseer):

or are you looking for help on something else?

OpenStudy (anonymous):

i've tried the answer wolfram gave me but it doesn't seem to be correct. It's an online assignment. I'm guessing it has something to do with the 'non-trivial solution' part of the question.

OpenStudy (asnaseer):

so you tried this:\[\frac{10^{x + y}}{\ln(10)}+constant\]

OpenStudy (anonymous):

yeah i did

OpenStudy (asnaseer):

well I can't help you then - maybe it is a quirk of the online system? it may require you to enter the answer in a very specific manner.

OpenStudy (asnaseer):

what wolfram gave IS a non-trivial solution - see here for the definition of what non-trivial is: http://www.goiit.com/posts/list/algebra-what-are-trivial-and-non-trivial-solutions-932014.htm;jsessionid=E44E4747AE2931ABF868DDF9ECEE1545.node1#1002744

OpenStudy (asnaseer):

you could try entering the answer WITHOUT the constant - i.e.:\[\frac{10^{x + y}}{\ln(10)}\]that's the only other suggestion I can think of.

OpenStudy (anonymous):

thanks

OpenStudy (asnaseer):

yw